OpenStudy (anonymous):

Solve for x: 4x + 2 = x – 7

OpenStudy (anonymous):

4x + 2 = x - 7 subtract the x to the other side then youre left with 3x + 2 = -7 then you add 2 then you have 3x = -9 then divide by 3 and youre answer is x= -3

OpenStudy (anonymous):

hope it helps!

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

you're welcome(: